

********************************************************************
UARS WEEKLY STATUS REPORT - 13 OCT 2000
********************************************************************
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is not affected by the larger than normal pointing errors
since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
However, level2 processing is halted due to deficiencies in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem and plan to have
a patch by the end of October.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for predicted periods of UARS telemetry.
We are preparing to begin to reprocess the SUSIM UV spectral
irradiances to data version V21 in the next few weeks.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.
The UARS calendar for year 2001 is now available on the UARS
homepage, under "What's New" section. (http:/umpgal.gsfc.nasa.gov)
***********
CDHF
***********
HIGHLIGHTS
UOAS OA_SSPP_ATT
- Need to debug additional UOAS OA_SSPP_ATT gap scenarios:
2) gap at end of day, time change, no gap at begin of next day.
3) tests run in RSS need to be run in PDS mode and possibly Menu mode.
- Need to code comparable fix in OA_SAT_ATT and test in debug.
- Plan to have patch by end of month.
DAAC DATA:
- MLS Version 5 data is being previewed and loaded into the DAAC Transfer
Tables. When all of the Version 5 has been loaded into the DAAC tables,
various fields will need updated before DAAC (James Johnson) can reveal/
make available to the User Community.
- Waiting response from WINDII that Version 11 Quality fields have been
updated.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE,
HRDI, MLS, SUSIM, Solstice, and WINDII.
- SUSIM v7.2 software representative Ed Esfandiari notified parallel
testing completed. *SUSIM responded with concerns regarding the output
product which is being investigated.
- HRDI v4.4 software representative Mr. Jim Raines was notified parallel
testing completed. *HRDI responded and is still in process of validating
the output product.
- GPI V4.05 tested successfully against the HRDI V4.4 Level 3AL data.
Frank Huang was requested to validate the GPI output product. *Frank
responded with questions and some concerns.
- SOLSTICE v4.62 software representative PI Barry Knapp has successfully
validated the parallel testing output.
- HALOE v4.82 software representative Mr. Lance Deaver was notified the
parallel test days completed. *HALOE has responded that they see no major
differences in the output product, but have 3 more species to check out
and will respond further this week. Small differences are possibly
attributed to interpolation routines.
- WINDII V5.11A software representative Mr. Brian Solheim requested parallel
processing on test days. Testing continues.
- July data ingested on the Rehost System for NMC and UKMO was successfully
previewed and loaded into the DAAC Transfer Tables. DAAC (James Johnson)
has been notified to test his transfers.
***********
FOT
***********
REPORTING PERIOD
This report covers 30 September 2000 (Orbit 49486, GMT Day 274) through 06
October 2000 (Orbit 49590, GMT Day 280).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 279/0532 49564 Off
HALOE 279/0953 49566 On
HALOE 280/1232 49584 Off
HALOE 280/1555 49586 On
The beta angle ranged between the angles of 44.1 degrees to 51.0 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
2 degrees on 23 October 2000 (DOY 297).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.0 degrees C. Additional battery performance data for each
battery on 29 September 2000 and 06 October 2000 are provided in Appendix D.
The clock error ranged from -12.7 to 11.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 01
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 25
¥ HALOE = 04
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/221 - 2287.496937
· Transponder B: 00/244 - 2287.499127
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/272 show the observatory data loss to be 298 hours, 29
minutes, 18.483 seconds (an increase of 2 minutes 18.252 seconds since the last
report period). This is a 0.39 percent data loss which equals a 99.61 percent
data capture for the mission.
The increase of 2 minutes 18.252 seconds occurred over two (2) days:
· 1 minute 20.672 seconds on Day 276 due to a data hit,
· 0 minutes 58.840 seconds on Day 278 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT powered off FHST-1 on 29 September 2000. The tracker will remain off
and UARS will fly on FSS and gyros alone. Should the attitude start to drift
the FOT has implemented a procedure to manually correct the attitude via an FDF
calculated attitude correction. UARS will continue to fly in this manner until
the TAM and FSS attitude control solution is put into place. Testing continues
daily to get this mode operational.
MEETINGS
The FOT attended a meeting to discuss UARS attitude control on 02 and 04
October 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for October 22, 2000 (DOY 297). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 Oct 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
49489 274/0525 00-525 NBTR NOOPS D274
49495 274/1250 00-526 OBC Clk Adj 20.6 to 0.0 msec/day
49497 274/1726 00-525 NBTR NOOPS D274
49497 274/1806 00-527 NBTR NOOPS D275
49512 275/1809 00-528 NBTR NOOPS D276
49520 276/0610 00-528 NBTR NOOPS D276
49526 276/1549 00-529 OBC GENASCII/ TABLE 51
49527 276/1811 00-530 NBTR NOOPS D277
49531 277/0052 00-531 OBC Clk Adj 0.0 to 20.6 msec/day
49535 277/0612 00-530 NBTR NOOPS D277
49540 277/1409 00-532 OBC GENASCII/TABLE 51
49543 277/1910 00-533 NBTR NOOPS D278
49544 277/2132 00-534 TAM2 GENASCII/TABLE 53
49549 278/0531 00-533 NBTR NOOPS D278
49555 278/1411 00-536 OBC GENASCII/TABLE 51
49556 278/1552 00-537 OBC GENASCII/TABLE 15
49557 278/1814 00-535 NBTR NOOPS D279
49564 279/0532 00-535 NBTR NOOPS D279
49564 279/0532 00-538 OBC Clk Adj 20.6 to 0.0 msec/day
49564 279/0532 00-539 HALOE POWER OFF/ TBAT HI
49566 279/0953 00-540 HALOE POWER ON
49570 279/1413 00-541 OBC GENASCII/TABLE 51
49572 279/1733 00-535 NBTR NOOPS D279
49572 279/1816 00-542 NBTR NOOPS D280
49579 280/0534 00-542 NBTR NOOPS D280
49584 280/1232 00-543 HALOE POWER OFF
49585 280/1415 00-544 OBC GYRO PREDICTS
49586 280/1555 00-545 HALOE POWER ON
49587 280/1817 00-546 NBTR NOOPS D281
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 Oct 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HATBAT Y-HI BOLOMETER AREA TMP 49581-82
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 Oct 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
49495 274/1250 (30 September) from 20.6 to 0.00 msec/day
49531 277/0052 (03 October) from 0.00 to 20.6 msec/day
49564 279/0532 (05 October) from 20.6 to 0.00 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 Oct 2000
29 Sept 2000 - GMT Day 273 Beta = 41.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-78.4 +5.47/+4.51 +27.2/-15.2 25.1 1.030 19.7
3 +16.8/-44.8 +4.51/+3.23 +27.2/-15.2 25.1 1.020 19.3
06 Oct 2000 - GMT Day 280 Beta = 51.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/ MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/+0.0 14.2 0. 000 0.0
2 +16.8/-56.0 +4.83/+3.23 +24.8/-15. 2 25.4 1.040 20.5
3 +11.2/-72.8 +3.87/+2.29 +25.2/-15. 2 24.8 1.040 20.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 Oct 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-042 FOT 49534 277/0529 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-041 FOT see list Sept 00 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-042 FOT see list Oct 00 14170 Generic Late Acq.
00-040 FOT 48854 231/2111 22831 Data Dropout
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
SEVNTFW00/40